what is (2)\(3)y - 6=2
Answer by Flannery(124)

I hope you mean 2/3y-6=2
2/3y=8 (add 6)
y=12 (divide by 2/3)(that is take the reciprocal and multiply)
Here's the check
2/3(12)-6=2
8-6=2
2=2
Q.E.D.